OpenStudy (anonymous):

Given the ordered pair (3, y), what value of y would make the ordered pair a solution of the equation 4x − 2y = 24 a 12 b 6 c-6 d-8

OpenStudy (kaiz122):

we know that x=3 substitute x to the equation

OpenStudy (kaiz122):

then you can simply solve for y, can you do it now?

OpenStudy (anonymous):

can you show me what your talking about? i dont know how to do this

OpenStudy (kaiz122):

since x=3 we can have 4(3)-2y=24 12-2y=24 2y=12-24

OpenStudy (kaiz122):

then solve for y

OpenStudy (anonymous):

-6

OpenStudy (kaiz122):

that's it